OpenStudy (anonymous):

in triangle ABC, measurement of Angle C=90 and line AC is congruent to line BC. If AB=12, find AC

OpenStudy (johnweldon1993):

Congruent means the "same" So if you have an angle of 90 you know you have a right triangle....but since we also know that 2 lines are the same...we know it is a 45 45 90 triangle  In a 45 45 90 triangle..the hypotenuse (the longest side so AB for you) would be the result of one of the smaller legs times √2  So we need to find out...WHAT times the √2 = 12 \[\large x \times √2 = 12\] \[\large x = \frac{12}{√2}\] \[\large x = \frac{12√2}{√2√2}\] \[\large x = \frac{12√2}{2}\] \[\large x = 6√2\] That is what you AC equals

OpenStudy (anonymous):

oh I was not sure if I was going down the right path thank you

OpenStudy (johnweldon1993):

No problem!